OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 13x - 6y = 20 7x + 4y = 18 2 1 -2 -1

OpenStudy (anonymous):

set one of the equations to be x= and then plug it into the other equation

OpenStudy (anonymous):

in this equation u say screw the Y's! 13x-6=20 so x is = 2 then u look at the second problem so u know x=2 7x+4y=18 7*2=14 so its 14+4y=18 y = 1 so your answers are x =2 y=1